OpenStudy (anonymous):

Which mathematical sentence expresses the information below? Jake has 32 quarters in his pocket, and Molly has q quarters in her purse. They have a total of 57 quarters

OpenStudy (anonymous):

T is for total. T=32 + q 57 - 32 is 25 57=32+25